Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point?
Answer by sdmmadam@yahoo.com(530)

Draw a rough sketch as directed please.
I don't know to draw graphs with in this box. My graphs from my applications refuse to get copied into this box
Start from a point A
Go vertically up to a point B
Mark this length AB = 65
Go horizontal from B to a point C with length BC slightly longer than AB.
Mark this length BC= 75
You started your journey at A. Went up north to B. Went east from B to C.
If now you join A and C, then the length CA gives the distance from your destination C to the starting point A.
If you look at your diagram, it is a right angled triangle with
angle B = 90 degrees. AC is the hypotenuse.
So by Pythog. theorem
AC^2 = AB^2 + BC^2
=(65)^2 + (75)^2
= 4225 +5625
=9850
Then AC = sqrt(9850)which is approximately 99.25 kilometers
Answer: The required distance is approximately 99.25 kilometers